A skier at the top of a hill has a velocity of 6m/s and when she reaches the bottom her velocity is 13 m/s. Ignore friction, and find how tall the hill was.
Answer #1

we can use the conservation of energy to solve this problem,

The sum of Kinetic energy =Sum of the potential energy

i.e

(1/2)mvi2+(1/2)mvf2=mgh

Once it reaches ground the final height =0 or the final potential energy=0

vi=inital velocity

vf=Final velocity

(1/2)m62+(1/2)m132=mgh

m is common in both sides, so let us cancel in both sides

So,1/2(36+169)=gh

102.5=gh

h=10.459 m

Height of the hill is 10.46 m